OpenStudy (anonymous):

solve by the elimination method: 4x + 5y = 3 8x + 10y = 6 what is the solution of the system? a. the solution is --- b. there are infinitely many solutions c. no solution

OpenStudy (anonymous):

b

OpenStudy (anonymous):

4x + 5y = 3 8x + 10y = 6 Multiply equation a by -2 as multiplying by a constant will not change the equation. Then you have: -8x - 10y = -6 8x + 10y = 6 If we add these together, we have: 0x + 0y = 0 So there is no definite answer and there are an infinite number of solutions as x and y can take on any value in the coordinate system. If this helps, please choose good answer. Visit my website at http://www.tutorsean.net for free math tutoring videos.